# Kalahi: A handcrafted, grassroots cultural LLM evaluation suite for Filipino

### v1

@@ Abstract
Multilingual large language models (LLMs) today may not necessarily provide culturally appropriate and relevant responses to its Filipino users. We introduce Kalahi, a cultural LLM evaluation suite collaboratively created by native Filipino speakers. It is composed of 150 high-quality, handcrafted and nuanced prompts that test LLMs for generations that are relevant to shared Filipino cultural knowledge and values. Strong LLM performance in Kalahi indicates a model’s ability to generate responses similar to what an average Filipino would say or do in a given situation. We conducted experiments on LLMs with multilingual and Filipino language support. Results show that Kalahi, while trivial for Filipinos, is challenging for LLMs, with the best model answering only 46.0% of the questions correctly compared to native Filipino performance of 89.10%. Thus, Kalahi can be used to accurately and reliably evaluate Filipino cultural representation in LLMs.
@@
Jann Railey Montalan1, 2, Jian Gang Ngui1,2, Wei Qi Leong1,2, Yosephine Susanto1,2, Hamsawardhini Rengarajan1,2, William Chandra Tjhi1,2, Alham Fikri Aji3,4 1AI Singapore, 2National University of Singapore, 3MBZUAI, 4Monash Indonesia Correspondence: railey@aisingapore.org
@@ 1 Introduction
To bridge this gap, we present Kalahi, [^1] Kultural na Analisis ng LLMs sa Ating PagpapaHalaga at Identidad (Cultural Analysis of LLMs on Our Values and Identity). The Filipino word kalahi (noun) means ‘someone from the same people, race, or origin’. This reflects our core belief that cultural evaluations should aim to test if an LLM can respond as if it ‘belongs’ or ‘acts like’ a member of a particular group of people or culture. a high-quality, manually-crafted cultural dataset to determine LLMs’ abilities to provide relevant responses to culturally-specific situations that Filipinos face in their day-to-day lives.
@@ 1.1 Contributions
    We present Kalahi, [^2] We will release our dataset and evaluation codebase on HuggingFace and GitHub soon. an evaluation suite with high-quality, handcrafted prompts that test the ability of LLMs to generate responses relevant to Filipino culture in terms of shared knowledge and ethics.
@@ 3.1 Manual Dataset Construction
We started by sourcing pertinent information from Google Trends, including most frequently searched terms, news, and YouTube queries in the Philippines from 2018 to 2023. The most popular search queries made in the Philippines were generally information-seeking (e.g. news on COVID), for practical tasks (e.g. English-Filipino translation), and for entertainment (e.g. song lyrics).
@@
The first iteration of prompt-response pairs was expanded to include a total of 150 prompts, each with accompanying personal situation variations. Each prompt has three to five relevant and irrelevant responses, with only one of the relevant responses being labeled the ‘best response’. [^5] We provide additional examples in Appendix C.
@@
Prompt-response categories. We present the cultural topics covered in Kalahi (see Table 3). Recall that we did not restrict ourselves to a predetermined set of topics, though we took some topics that were found to be important as a starting point for the FGDs. We highlight below the motivation behind grouping certain topics together:
@@ 3.2 Dataset Validation
We recruited three native Filipino speakers who were not involved in the development of Kalahi to validate the constructed dataset. We evaluate the validators on the MC1 task (see Section 4.2). These validators were shown the 150 prompts from Kalahi and best and irrelevant responses in a randomized order. They were tasked to choose the response that would most closely mirror the choice that an average Filipino would make given a particular situation as their ‘strategy of action’. It is important to remember that the irrelevant responses could sound innocuous, logical, or reasonable in the context of other cultures, but crucially they are rendered irrelevant in Filipino culture (i.e. such responses would not be strategies of actions adopted by the average Filipino). The three native speakers attempted all 150 prompts and these validator answers were then used as the human baseline for our experiments
@@ 4.1 Human baseline
On average, our Filipino validators scored 89.1% on Kalahi, which we refer to as our human baseline. We calculated inter-rater agreement, which yielded a Cohen’s kappa of 0.761 and a Krippendorf’s alpha of 0.762, indicating substantial agreement. While Kalahi was created based on consensus among native Filipinos, individual idiosyncrasies, such as personal values and beliefs, were expected to inherently influence their individual choices, such that the participants’ choices may not necessarily align with the shared Filipino cultural values and beliefs. This can be observed in the example in Appendix D.
@@
4.2 Model Evaluation
Experiments. We evaluate a total of 9 LLMs to compute baselines for Kalahi. The first group of LLMs are those that explicitly claim to support Filipino (Tagalog), which we assume means that the models were instruction-tuned on Filipino instructions: Aya 23 8B Aryabumi et al. (2024), Qwen 2 7B Instruct Yang et al. (2024), Sailor 7B Chat Dou et al. (2024), and SeaLLMs 3 7B Chat Zhang et al. (2024). The second group are composed of LLMs that have claimed to demonstrate multilingual capabilities, but do not claim to be specifically instruction-tuned on Filipino instructions: BLOOMZ 7B1 BigScience Workshop et al. (2023), Falcon 7B Instruct Almazrouei et al. (2023), Gemma 2 9B Instruct Gemma Team et al. (2024), Llama 3.1 8B Instruct Dubey et al. (2024), and SEA-LION 2.1 8B Instruct.
@@ 4.2 Model Evaluation
Table 5: Model performance on Kalahi for both settings (see Appendix E for more results).
Multiple-choice. In this setting, a model is evaluated on a multiple-choice question. The choices for each question refer to relevant and irrelevant responses. We compute the log-probability completion of each reference response given a question, normalized by byte length. Two scores are calculated:
@@ 4.3 Interpretation of Results
Models that claim to support Filipino do provide more culturally relevant responses. Results show that LLMs that claim Filipino langauge support (presumably meaning they have seen more Filipino tokens in training) generally perform better on Kalahi. Table 5 shows that these LLMs also had strong performance in open-ended generation. A manual review of generated completions from these models showed that their responses were in Filipino and provided actionable suggestions that the user could take. The following example is illustrative:
@@
Gemma 2 9B: “Ibigay ang sapat na pahinga… hindi pinipilit na maglaro… Panatilihing hydrated: Ibigay sa kanya ang maraming tubig… Ibigay sa kanya malusog at malambot na pagkain na madaling lunukin…” (“Give enough rest… don’t force them to play… Stay hydrated: drink lots of water… Give nutritious and soft foods that are easy to swallow…”)
@@
This insight is also supported by the fact that in the multiple-choice setting, LLMs with Filipino language support performed better. That is, the results seem to suggest that that models with Filipino language support also have stronger representations of the relevant aspects of Filipino culture. However, it should still be noted that while all of the models significantly outperformed random guessing, none of them performed as well as a native Filipino.
@@
For the multiple-choice setting, literature suggests that models may not understand the task of selecting options Khatun and Brown (2024), are sensitive to the order in which options are presented Leong et al. (2023); Pezeshkpour and Hruschka (2023); Wang et al. (2024a), and first token probabilities may not be reflective of what the model would generate Li et al. (2024b); Lyu et al. (2024); Wang et al. (2024b). We mitigate these problems by not presenting the models with a list of options and requesting for a single letter as a response, and instead measuring the probability of each option separately.
@@
Although Kalahi is fully-automated, we also recommend manually validating the model’s output as an additional verification process to further increase reliability of the results. Hallucinations and culturally inappropriate responses are not automatically detected in Kalahi, but it is notable that there were instances flagged out by our native speakers, such as the following example:
@@
Llama 3.1 8B: “Dalhin ang iyong anak sa ospital upang mapag-aralan ng mga doktor ang kanyang kalagayan…” (“Bring your child to the hospital so that the doctors can study their condition…”)
@@ 5 Conclusion
Our findings show that multilingual LLMs and even those that have Filipino language support significantly underperform compared to native Filipinos on Kalahi. This demonstrates that Kalahi is a challenging benchmark for evaluating Filipino cultural representation in LLMs.
@@
Limitations. The Filipino culture in this study refers to cultural values acquired by Filipino speakers who were born and grew up in or at least spent most of their lives in the Philippines. Kalahi is the result of the consensus views of the involved native Filipino speakers. Individuals who have had different upbringings may have different perspectives on Filipino culture mentioned in this paper, such that the consensus view arrived at in this study does not fully represent the opinions of all Filipino individuals. Additionally, while Kalahi is designed to accurately represent Filipino culture, it is not intended to encompass all possible aspects of Filipino culture.
@@ Appendix A Data construction guidelines
Given the subjectiveness of ‘culture’, it is infeasible to adopt a normative stance. We instead adopt a more collaborative approach that involves native speakers from the respective communities to help inform the data collection process. This set of data construction guidelines [^7] The guidelines have been reviewed and approved by an Institutional Review Board (IRB). is intended to detail a methodology for researchers who are looking to collect data from the community in a principled manner.
@@
Appendix B Focus group discussion participant demographics
@@
Table 6: Demographics of focus group discussion (FGD) participants.
Appendix C Additional prompt-response pair examples
@@
Appendix D Disagreement between personal and cultural values and beliefs of Filipino validators
@@
Appendix E Open-ended generation model performance
@@
Appendix F Ablation study: model performance on prompts without enriching contexts
@@
Table 11 shows that models’ performances are not consistently affected by the removal of ‘User’. For instance, while we observe that Aya 23 8B’s performance on the MC1 task improved, Gemma 2 9B Instruct’s performance deteriorated. Interestingly, SeaLLMs 3 7B Chat’s performance was unaffected. The results in Table 12 also show that models’ performances are not consistently affected. We hypothesize that the inconsistency is an indication that the models are easily perturbed, especially considering that they generally do not perform on Kalahi regardless.
@@
Interestingly, MC1 and MC2 performance in Table 13 generally shows that all models’ performances deteriorated on Kalahi unenriched prompts. However, performance on the open-ended generation setting in Table 14 shows that Llama 3.1 8B Instruct’s performance improved (and drastically so, for some of the metrics) on Kalahi unenriched prompts. [^8] Note that there are only 84 Kalahi unenriched prompts in comparison to the full 150 Kalahi fully-enriched prompts.

### v2

@@ Abstract
Multilingual large language models (LLMs) today may not necessarily provide culturally appropriate and relevant responses to its Filipino users. We introduce Kalahi, a cultural LLM evaluation suite that is part of SEA-HELM. It was collaboratively created by native Filipino speakers, and is composed of 150 high-quality, handcrafted and nuanced prompts that test LLMs for generations that are relevant to shared Filipino cultural knowledge and values. Strong LLM performance in Kalahi indicates a model’s ability to generate responses similar to what an average Filipino would say or do in a given situation. We conducted experiments on LLMs with multilingual and Filipino language support. Results show that Kalahi, while trivial for Filipinos, is challenging for LLMs, with the best model answering only 46.0% of the questions correctly compared to native Filipino performance of 89.10%. Thus, Kalahi can be used to accurately and reliably evaluate Filipino cultural representation in LLMs.
@@
Jann Railey Montalan1,2, Jian Gang Ngui1,2, Wei Qi Leong1,2, Yosephine Susanto1,2, Hamsawardhini Rengarajan1,2, Alham Fikri Aji3,4, William Chandra Tjhi1,2 1AI Singapore, 2National University of Singapore, 3MBZUAI, 4Monash Indonesia Correspondence: railey@aisingapore.org
@@ 1 Introduction
To bridge this gap, we present Kalahi, [^1] Kultural na Analisis ng LLMs sa Ating PagpapaHalaga at Identidad (Cultural Analysis of LLMs on Our Values and Identity). The Filipino word kalahi (noun) means ‘someone from the same people, race, or origin’. This reflects our core belief that cultural evaluations should aim to test if an LLM can respond as if it ‘belongs’ or ‘acts like’ a member of a particular group of people or culture. a high-quality, manually-crafted cultural dataset that is part of SEA-HELM [^2] https://leaderboard.sea-lion.ai/ and designed to determine LLMs’ abilities to provide relevant responses to culturally-specific situations that Filipinos face in their day-to-day lives.
@@ 1.1 Contributions
    We present Kalahi, an evaluation suite [^3] https://github.com/aisingapore/kalahi with high-quality, handcrafted prompts [^4] https://huggingface.co/datasets/aisingapore/kalahi that test the ability of LLMs to generate responses relevant to Filipino culture in terms of shared knowledge and ethics.
@@ 3.1 Manual Dataset Construction
We started by sourcing pertinent information from Google Trends, including most frequently searched terms, news, and YouTube queries in the Philippines from 2018 to 2023. The most popular search queries made in the Philippines were generally for information (e.g. news on COVID), practical tasks (e.g. English-Filipino translation), and entertainment (e.g. song lyrics).
@@
The first iteration of prompt-response pairs was expanded to include a total of 150 prompts, each with accompanying personal situation variations. Each prompt has three to five relevant and irrelevant responses, with only one of the relevant responses being labeled the ‘best response’. [^7] We provide additional examples in Appendix D.
@@
Prompt-response categories. We present the cultural topics covered in Kalahi (see Table 3). Recall that we did not restrict ourselves to a predetermined set of topics, though we took some topics that were found to be important as a starting point for the FGDs. Appendix C discusses the motivation behind grouping certain topics together.
@@ 3.2 Dataset Validation
We recruited three native Filipino speakers who were not involved in the development of Kalahi to validate the constructed dataset. We evaluate the validators on the MC1 task (see Section 4.2). These validators were shown the 150 prompts from Kalahi and best and irrelevant responses in a randomized order. They were tasked to choose the response that would most closely mirror the choice that an average Filipino would make given a particular situation as their ‘strategy of action’. It is important to remember that the irrelevant responses could sound innocuous, logical, or reasonable in the context of other cultures, but crucially they are rendered irrelevant in Filipino culture (i.e. such responses would not be strategies of actions adopted by the average Filipino). The three native speakers attempted all 150 prompts and these validator answers were then used as the human baseline for our experiments.
@@ 4.1 Human baseline
On average, our Filipino validators scored 89.1% on Kalahi, which we refer to as our human baseline. [^8] An interesting avenue for future work would be to have considerably more Filipinos attempt Kalahi to set a stronger human baseline as well as to mitigate personal biases. We calculated inter-rater agreement, which yielded a Cohen’s kappa of 0.761 and a Krippendorf’s alpha of 0.762, indicating substantial agreement. While Kalahi was created based on consensus among native Filipinos, individual idiosyncrasies, such as personal values and beliefs, were expected to inherently influence their individual choices, such that the participants’ choices may not necessarily align with the shared Filipino cultural values and beliefs. This can be observed in the example in Appendix E.
@@ 4.2 Model Evaluation
Experiments. We evaluate a total of 9 LLMs to compute baselines for Kalahi. The first group of LLMs explicitly claim to support Filipino (Tagalog), which we assume means that the models were instruction-tuned on Filipino instructions: Aya 23 8B Aryabumi et al. (2024), Qwen 2 7B Instruct Yang et al. (2024), Sailor 7B Chat Dou et al. (2024), and SeaLLMs 3 7B Chat Zhang et al. (2024). The second group include LLMs that have claimed to demonstrate multilingual capabilities, but do not claim to be specifically instruction-tuned on Filipino instructions: BLOOMZ 7B1 BigScience Workshop et al. (2023), Falcon 7B Instruct Almazrouei et al. (2023), Gemma 2 9B Instruct Gemma Team et al. (2024), Llama 3.1 8B Instruct Dubey et al. (2024), and SEA-LION 2.1 8B Instruct.
@@
Table 5: Model performance on Kalahi for both settings (see Appendix G for more results).
Multiple-choice. In this setting, a model is evaluated on a multiple-choice question. The choices for each question refer to relevant and irrelevant responses. We compute the log-probability completion of each reference response given a question, normalized by byte length. Two scores [^9] Appendix F illustrates how MC1 and MC2 are calculated. are calculated:
@@ 4.3 Interpretation of Results
Models that claim to support Filipino do provide more culturally relevant responses. Results show that LLMs that claim Filipino langauge support (presumably meaning they have seen more Filipino tokens in training) generally perform better on Kalahi. Table 5 shows that these LLMs also had strong performance in open-ended generation. A manual review of generated completions (see Appendix I) from these models showed that their responses were in Filipino and provided actionable suggestions that the user could take. The following example is illustrative:
@@
Gemma 2 9B Instruct: “Ibigay ang sapat na pahinga… hindi pinipilit na maglaro… Panatilihing hydrated: Ibigay sa kanya ang maraming tubig… Ibigay sa kanya malusog at malambot na pagkain na madaling lunukin…” (“Give enough rest… don’t force them to play… Stay hydrated: drink lots of water… Give nutritious and soft foods that are easy to swallow…”)
@@
This insight is also supported by the fact that in the multiple-choice setting, LLMs with Filipino language support performed better. That is, the results seem to suggest that that models with Filipino language support also have stronger representations of the relevant aspects of Filipino culture. However, it should still be noted that while all of the models outperformed the random baseline, none of them performed as well as a native Filipino.
@@
For the multiple-choice setting, literature suggests that models may not understand the task of selecting options Khatun and Brown (2024), are sensitive to the order in which options are presented Leong et al. (2023); Pezeshkpour and Hruschka (2023); Wang et al. (2024a), and first token probabilities may not be reflective of what the model would generate Li et al. (2024b); Lyu et al. (2024); Wang et al. (2024b). We mitigate these problems by not presenting the models with a list of options and requesting for a single letter as a response, and instead measuring the log probability of each option as a model completion separately.
@@
Although Kalahi is fully-automated, we also recommend manually validating the model’s output as an additional verification process to further increase reliability of the results. [^11] We conducted human evaluations on subsets of model generations and reported preliminary findings in Appendix I. Hallucinations and culturally inappropriate responses are not automatically detected in Kalahi, but it is notable that there were instances flagged out by our native speakers, such as the following example:
@@
Llama 3.1 8B Instruct: “Dalhin ang iyong anak sa ospital upang mapag-aralan ng mga doktor ang kanyang kalagayan…” (“Bring your child to the hospital so that the doctors can study their condition…”)
@@ 5 Conclusion
Our findings show that multilingual LLMs and even those that have Filipino language support still underperform compared to the native Filipino baseline on Kalahi. This demonstrates that Kalahi is a challenging benchmark for evaluating Filipino cultural representation in LLMs.
@@
We also encourage researchers to conduct surveys with larger groups of native speakers, in collaboration with cultural experts, linguists, sociologists, and anthropologists in order to collect more culturally representative data.
Limitations. While Kalahi is the result of the consensus views of the involved native Filipino speakers, the Filipino culture in this study refers only to cultural values acquired by Filipino speakers who were born and grew up in or at least spent most of their lives in Metro Manila. Individuals who have had different upbringings may have different perspectives on Filipino culture, such that the consensus view arrived at in this study does not fully represent the opinions of all Filipino individuals. Additionally, while Kalahi is designed to accurately represent Filipino culture, it is not intended to encompass all possible aspects of Filipino culture.
@@ Appendix A Data construction guidelines
Given the subjectiveness of ‘culture’, it is infeasible to adopt a normative stance. We instead adopt a more collaborative approach that involves native speakers from the respective communities to help inform the data collection process. This set of data construction guidelines [^12] The guidelines have been reviewed and approved by an Institutional Review Board (NUS-IRB-2024-617). is intended to detail a methodology for researchers who are looking to collect data from the community in a principled manner.
@@
Appendix B Demographics of focus group discussion participants
@@
Table 6: Demographics of focus group discussion (FGD) participants. All participants spent a majority of their lives and currently live in Metro Manila.
Appendix C Grouping of cultural topics
The motivation for the categorization of some of the cultural prompts in Kalahi are as follows:
Appendix D Additional prompt-response pair examples
@@
Appendix E Disagreement between personal and cultural values and beliefs of Filipino validators
@@
Appendix F Illustration of log-probability calculation for MC1 and MC2
The implementations of the MC1 and MC2 scores are derived from TruthfulQA, Lin et al. (2022). While the MC1 and MC2 scores in TruthfulQA measure the ‘truthfulness’ of model responses, we reframe these scores as measurements of cultural relevance of model responses in this study.
It should be noted that for the MC1 task, as long as the log-probability for the ‘best response’ label turns out to be the highest, the model will receive a score of 1. However, such a scoring method obscures the differences in log-probabilities assigned to the other labels.
The MC2 task addresses this by providing a value that indicates whether the summed log-probabilities of the relevant responses are higher or lower than that of the irrelevant responses. Indeed, given the scores of the models in Table 5, it seems to indicate that the differences in log-probabilities of relevant and irrelevant responses are potentially insignificant.
Figure 2: Calculation for the MC1 metric.
Figure 3: Calculation for the MC2 metric.
Appendix G Open-ended generation model performance
@@
Appendix H Ablation study: model performance on prompts without enriching contexts
@@
Table 11 shows that models’ performances are not consistently affected by the removal of ‘User’. For instance, while we observe that Aya 23 8B’s performance on the MC1 task improved, Gemma 2 9B Instruct’s performance deteriorated. Interestingly, SeaLLMs 3 7B Chat’s performance was unaffected. The results in Table 12 also show that models’ performances are not consistently affected. We hypothesize that the inconsistency is an indication that the models are easily perturbed, especially considering that they generally do not perform well on Kalahi regardless.
@@
Interestingly, MC1 and MC2 performance in Table 13 generally shows that all models’ performances deteriorated on Kalahi unenriched prompts. However, performance on the open-ended generation setting in Table 14 shows that Llama 3.1 8B Instruct’s performance improved (and drastically so, for some of the metrics) on Kalahi unenriched prompts.
@@
Appendix I Human evaluation of model open-ended generation
To further determine if the evaluated LLMs truly provide relevant responses under Kalahi, we conduct human evaluations to determine the helpfulness and harmlessness of the models’ generations. Four LLMs were evaluated: two models with Filipino language support (Qwen 2 7B Instruct and SeaLLMs 3 7B Chat), and two models without dedicated Filipino instruction tuning (Gemma 2 9B Instruct and Llama 3.1 8B Instruct). The model responses to 60 randomly-selected prompts, totaling to 240 unique responses, were evaluated. There were two groups composed of three native Filipino speakers each (for a total of six native speakers). Each group evaluated 120 of the 240 responses. The criteria for evaluation are as follows:
  1. 1.
    Factuality (Fac): The response does not contain any factual errors.
  2. 2.
    Grammaticality (Gra): The response does not contain any grammatical errors.
  3. 3.
    Spelling Correctness (Spe): The response does not contain any spelling errors.
  4. 4.
    Coherence (Coh): The response is relevant to the prompt and is not nonsensical or contains hallucinations.
  5. 5.
    Cultural Actionability (Cac): The response contains strategies of action that can be executed within the shared morals, restrictions, and preferences of the culture.
  6. 6.
    Cultural Sensitivity and Appropriateness (Csa): The response contains strategies of action that are not offensive within the culture.
  7. 7.
    Legality (Leg): The response contains strategies of action that are not illegal within the culture.
The results of the human evaluation based on the seven criteria are presented in Tables 16 and 17. For each criteria, we report the number of times that at least a majority (2/3) of the evaluators agreed that the model response demonstrated the criteria in question.
Model Fac Gra Spe Coh Cac Csa Leg
Models with Filipino language support
Qwen 2 7B Instruct 0.2500 0.4333 0.8333 0.3667 0.2500 0.9833 1.0000
SeaLLMs 3 7B Chat 0.5167 0.6000 1.0000 0.5500 0.3833 0.9500 0.9833
Multilingual models without dedicated Filipino instruction tuning
Gemma 2 9B Instruct 0.9333 0.9000 0.9833 0.9833 0.7500 0.9833 1.000
Llama 3.1 8B Instruct 0.5000 0.5667 0.9333 0.6500 0.5667 0.9667 1.000
Table 16: Human evaluation of factuality (Fac), grammaticality (Gra), spelling correctness (Spe), coherence (Coh), cultural actionability (Cac), cultural sensitivity and appropriateness (Csa), and legality (Leg) of model responses on Kalahi.
Model Fac Gra Spe Coh Cac Csa Leg
Models with Filipino language support
Qwen 2 7B Instruct 15/60 26/60 50/60 22/60 15/60 59/60 60/60
SeaLLMs 3 7B Chat 31/60 36/60 60/60 33/60 23/60 57/60 59/60
Multilingual models without dedicated Filipino instruction tuning
Gemma 2 9B Instruct 56/60 54/60 59/60 59/60 45/60 59/60 60/60
Llama 3.1 8B Instruct 30/60 34/60 56/60 39/60 34/60 58/60 60/60
Table 17: Portion of responses that demonstrated the criteria according to a majority (2/3) of the human evaluators.
I.1 Hallucination may lead to factual errors and incoherence, but not vice versa
Since hallucinations are not always easy to detect, but do lead to factual errors as well as incoherence in model generations, we decided to ascertain the factuality (Fac) and coherence (Coh) of model responses instead. By having these two criteria, we are also able to detect generations that are ultimately non-factual or incoherent but are not due to hallucinations.
In the model responses found in Tables 18 and 19, the responses simply do not discuss the relevant targeted fact, hence coming under the criterion of (lack of) factuality. Furthermore, in Table 19, the fact that the model’s generation is gibberish is also not attributable to any hallucination, and instead comes under the criterion of (lack of) coherence.
Model Response Translated Response
Gemma 2 9B Instruct Ang pagtalon sa pagpatak ng Bagong Taon ay isang tradisyon na ginagawa ng mga bata (at minsan ng mga matatanda!) sa buong mundo. Mayroong ilang mga dahilan kung bakit ito ginagawa… Pag-alis ng bad luck… Pag-asa para sa isang mas magandang taon… Pagdiriwang at kasiyahan… Jumping exactly on the arrival of the New Year is a tradition done by children (and sometimes adults!) all over the world. There are several reasons why this is done… Remove bad luck… Hope for a good year… Celebration and joy…
Table 18: A model response demonstrating an error in factuality (Fac) for the prompt: “Magbabagong taon na. Nagpaplano ang mga bata na tumatalon sa pagpatak ng Bagong Taon. Bakit ginagawa ito?” (“They new year is about to arrive. The children are planning to jump at the stroke of midnight of the New Year. Why is this done?”).
The belief is that one would grow taller if they jump exactly as the clock strikes midnight on New Year’s day. The response, however, does not mention anything about growing in height, instead mentioning other New Year’s beliefs such as bringing luck and happiness into the coming year.
Model Response Translated Response
Qwen 2 7B Instruct Ang kasalukuyang sitwasyon ninyo ay nagbibigay-daan sa inyo na magpa-rehat [sic] at magpababa ng tansihan. Ang paglamay o paglalayag ng katawan ay isang paraan ng pamamahala ng stress at pagkakaroon ng pag-asa sa pagbabago ng kahirapan at paniniwala sa magandang kalagitnaan. Sa panahon ng pagpapagpag, ang inyong katawan ay natututunan na magpasok sa isang estado ng ‘relaksasyon’ o ‘rest and digest’ mode, na mayroong positibong epekto sa iyong kakayahang maging mas mataas ang imunohimikal, maging mas malaki ang apoy sa ating mga organo… Your current situation gives way for you to rehat [sic] and lower your lamentation. Attending a wake or the exploration of your body is a way to handle stress and hope for change in poverty and believe in a beautiful middle. During the time of pagpag, your body learns to enter a state of ‘relaxation’ and ‘rest and digest’ mode, with positive effects in your ability to have better immunity, to have a stronger flame in your organs…
Table 19: A model response demonstrating an error in coherence (Coh) for the prompt: “Kakagaling ko lang mula sa lamay. Sabi ng magulang ko na ‘magpagpag’ muna ako bago umuwi sa bahay. Bakit ginagawa ito?” (“I just came from a wake. My parents said that I should ‘pagpag’ first before going home. Why is this done?”).
First, and most importantly, the response is mostly gibberish. Second, the response fails to address the concept of pagpag, which is the practice of going somewhere else first before heading home after a wake or funeral. Instead, it focuses on lamay, which is the Filipino word for wake. Third, the response hallucinates that lamay is a way to manage stress.
I.2 Human evaluation is necessary for understanding legality, cultural actionability, and cultural sensitivity and appropriateness
The concepts of legality (Leg), cultural actionability (Cac), and also cultural sensitivity and appropriateness (Csa) largely vary between cultures and countries. As most commericial LLMs like ChatGPT have been shown to demonstrate a Western, Industrialized, Rich, Educated, and Democratic (WIRED) bias Atari et al. (2023), using such models for automatic evaluation on cultural topics would be infeasible. Therefore, it is necessary for individuals belonging to the given culture to evaluate model generations on these criteria.
We note that almost all of the model generations demonstrated the criteria of legality (Leg) and cultural sensitivity and appropriateness (Csa), showing the effectiveness of their respective safety alignment approaches even within the Filipino language setting. There were some exceptions, such as the responses in Table 20 and 21, but these were far and few between.
Model Response Translated Response
SeaLLMs 3 7B Chat Magpautang ng Paspor [sic]: Sabihin sa kawanggawa ng isa na may pasaport at humingi ng pautang. Maaari kang magpakita ng mga alahas… Loan out your Paspor [sic]: Tell the charity of one who has a passport and ask for a loan. You can show jewelry…
Table 20: A model response demonstrating an error in legality (Leg) for the prompt: “Nasa hustong gulang na ako. Nais kong makakuha ng passport ng mabilisan. Anong maaari kong gawin?” (“I am an adult. I want to get a passport quickly. What can I do?”).
It is illegal to borrow or lend a passport to another person. Furthermore, the phrase “You can show jewelry…” could be understood as a recommendation to bribe someone (such as personnel at the passport agency) with jewelry, which is also illegal.
Model Response Translated Response
Qwen 2 7B Instruct Sa sitwasyon na ito, maaaring gawing pangako o salamat kay lola mo. Halimbawa: (1) T***ina, salamat sa ginhawa at pagmamasid… In this situation, you can make a promise or thank your grandmother. For example: (1) S** of a b***, thank you for the relief and regards…
Table 21: A model response demonstrating an error in cultural sensitivity and appropriateness (Csa) for the prompt: “Siyam na taong gulang ako. Nasa isang family reunion ako ngayon. Inabutan ako ng lola ko ng kanyang kamay. Anong maaari kong gawin?” (“I am nine years old. I am in a family reunion right now. My grandmother extended her hand to me. What should I do?”).
The response included an expletive that is offensive, especially towards an older relative to whom respect should be shown.
It is also notable that, on average, all of the models evaluated scored the weakest in cultural actionability (Cac) compared to all other evaluation criteria. Table 22 illustrates a model response that is culturally irrelevant by not providing actionable suggestions for the average Filipino. We propose that such divergence from the preferred strategies of actions that Filipino natives would take is an indicator that such models do not have strong representations of relevant aspects of Filipino culture.
Model Response Translated Response
Llama 3.1 8B Dalhin ang iyong anak sa ospital upang mapag-aralan ng mga doktor ang kanyang kalagayan… Tumawag sa mga serbisyong pangkalusugan sa iyong lugar upang makakuha ng tulong. Maaari silang magbigay ng mga gamot… Bring your child to the hospital so that the doctors can study their condition… Call health services in your area to get help. They may be able to give medicine…
Table 22: A model response demonstrating an error in cultural actionability (Cac) for the prompt: “May malubhang trangkaso ang walong taong gulang kong anak, at gusto kong gumaling na siya. Hindi ako mayaman. Anong maaari kong gawin?” (“My 8 year-old child has a severe flu, and I want them to get better. I am not rich. What can I do?”).
Because professional healthcare can be expensive and unaffordable in the Philippines, most Filipino households would not consider a doctor’s visit as a viable or top-of-mind option, preferring alternative treatments and home remedies instead. As such, this response is not culturally actionable by not providing such options.
I.3 Grammaticality, spelling correctness, and coherence serve as strong indicators for an LLM’s grasp of the Filipino language
What constituted a typographical error varied differently among human evaluators, demonstrating the subjectivity of what can be considered correct spelling. It has been noted in literature that speakers often use the incorrect characters to spell Filipino words, especially English loan words, because they are accustomed to the manner in which the English equivalents of such Filipino words are spelled Octaviano et al. (2016). As such, an incorrectly-spelled word may appear to be correctly-spelled to an evaluator even if it categorically demonstrates typographical errors. Therefore, it is recommended to use official guidelines for writing Filipino such the KWF Manwal sa Masinop na Pagsulat by Almario (2014) for evaluating typographical errors.
That said, the typographical errors that stand out in model generations, such as the one in Table 23, are not typical errors that native speakers would make, hence we assume that their presence is indicative that the model’s grasp of Filipino is still inadequate.
Model Response Translated Response
Qwen 2 7B Instruct Oo, ang ganoon ay malaking posibilidad na ang iyong plano ay magkakorrespon [incorrect spelling] sa iyong oras at pangunahing taktika… Yes, that is a big possibility that your plan will correspond to your time and primary tactic…
Table 23: A model response demonstrating an error in spelling correctness (Spe) for the prompt: “Nasa hustong gulang na ako. Pupunta ako sa mall sa Sabado. Gusto kong manamit ng pambahay lamang. Angkop ba ito?” (“I am an adult. I will go to the mall on Saturday. I want to wear only home clothes. Is this appropriate?”).
The response contains the word magkakorrespon, which is not a valid Filipino word. It seemingly reduplicates the letter ‘r’, similar to how the English verb correspond has two ‘r’s. In Filipino, this word should be spelled as magkokoresponda or magko-correspond (code-switched); the verb Filipino magtutugma (match) could also be used in lieu of this.
Another signal we find to be indicative is when models apply incorrect Filipino conjugations, which result in the overall generation being incoherent. Again, the errors such as those in Table 24 are not typical errors a native speaker would make since native speakers would have a strong grasp of Filipino conjugations and grammatical rules in general.
Model Response Translated Response
SeaLLMs 3 7B Chat … Sa kasong ito, kung ang iyong boss ay niyaya [incorrect conjugation of yaya] ang iyong pagkain [incorrect conjugation of kain] kasama ang mga katrabaho… … In this case, if your boss was invited [incorrect conjugation] (the act of) eating [incorrect conjugation] with your co-workers…
Table 24: A model response demonstrating an error in grammaticality (Gra) for the prompt: “Nasa hustong gulang na ako. Niyaya ako ng boss ko na lumabas kami kasama ang aming mga katrabaho para kumain sa weekend. Hindi ako malapit sa kanya. Angkop ba ito?” (“I am an adult. By boss invited me to go out to eat with my co-workers this weekend. I am not close to them. Is this appropriate?”).
First, the response uses the incorrect conjugation of the Filipino verb yaya (invite): the object-focus verb niyaya (i.e. the boss was invited) should be replaced with the actor-focus verb nagyaya (i.e. the boss invited). Second, the response uses the incorrect conjugation of the Filipino verb kain (eat): the nominalized verb pagkain (the act of eating) should be replaced with the infinitive form kumain (to eat).